OpenStudy (anonymous):

HELP PLZ!! 60% of a number is 300. What is 25% of the number? A. 45 B. 255 C. 720 D. 125

OpenStudy (anonymous):

60/100*x=300 so x= 500 and 25% =1/4 *500=125

OpenStudy (anonymous):

oh okie. So, it's D

OpenStudy (anonymous):

Wait, but how did u get 500?

OpenStudy (anonymous):

Let the number be : \(x\). 60% of that number = 300 That means: \[60 \% \; of \; x = 300 \\ \frac{60}{100} \times x = 300\]

OpenStudy (anonymous):

Can you solve for \(x\) now??

OpenStudy (anonymous):

That would b 500

OpenStudy (anonymous):

eta has simply typed my answer differently x=500 and a quarter of it is 125 answer D